# Boosted GFlowNets Improving Exploration via Sequential Learning

### v1

@@ 5.1 Synthetic Multimodal Environments
Training protocol. We organized the experiments into three settings: (i) a single GFlowNet trained for 10,000 epochs (SINGLE), (ii) a baseline trained for 3,000 epochs followed by a first booster trained for an additional 7,000 epochs (BGFN-2), and (iii) the same baseline and first booster up to 6,000 epochs, after which a second booster is introduced and trained for 4,000 epochs (BGFN-3). To control for stochasticity, all models were trained with identical random seeds across conditions.
Results. Our results on the synthetic environments show a clear and consistent pattern, visualized in Figures 1 and 2, and Table 1:
@@ 5.2 Antimicrobial Peptide (AMP) Discovery
We evaluate BGFNs on a real combinatorial design task: the generation of antimicrobial peptides (AMPs) trabucco2022designbenchbenchmarksdatadrivenoffline; pirtskhalava_dbaasp. Unlike prior work that fixes sequence length, here peptides have variable length from 1 to 10 amino acids. We train a Random Forest proxy to predict the probability of antimicrobial activity and define the GFlowNet reward as the logit of this probability. To avoid overweighting a small set of very-high-scoring sequences, we apply a 94% cutoff and clip the log-reward at zero for sequences above this threshold (i.e., all peptides with predicted activity $\geq 0.94$ share the same terminal reward). This choice preserves ranking signal below the threshold while mitigating oversampling to a few modes above it.
Training protocol (AMPs). All models are trained for 3,000 epochs. We activate Booster 1 at epoch 1,200 and Booster 2 at epoch 2,400, keeping earlier stages frozen thereafter. We evaluate both on-policy and off-policy training under estimator noise levels $\varepsilon\in\{0.1,0.2,0.3\}$. In addition, we compare two loss instantiations for incorporating past rewards: the flow-additive and the target-residual variants. As in the synthetic setup, we reuse the same random seeds across conditions to ensure that observed differences arise from the training regime rather than luck.
Results. Our results show the benefits in exploration and diversity obtained by BGFNs as shown in Figure 3 and Table 2. We observe that:
@@ 6 Conclusion, Limitations, and Broader Impact
Conclusion. We introduced Boosted GFlowNets (BGFNs), a framework for sequentially training ensembles of GFlowNets. Our results demonstrate that BGFNs are safe to use: new boosters can be added to improve exploration without degrading the performance of previously trained components. This property provides a principled way of enhancing exploration in challenging environments, where the topology of the state space creates strong biases towards easy-to-reach modes, even under off-policy training. Across synthetic multimodal benchmarks and antimicrobial peptide discovery, BGFNs consistently improved mode coverage and robustness, validating the framework as a reliable extension of standard GFlowNet training.
@@
Limitations. The main limitation of BGFNs is computational: each additional booster requires sampling from the backwards policy of all previously trained models in the ensemble. This increases training time on the order of $\mathcal{O}(KN)$, where $K$ denotes the number of trajectories sampled for a terminal node and $N$ the number of GFlowNets in the ensemble. An important exception is sequence environments with a unique backward path, in which the forward pass specifies the single valid backward trajectory; in practice, this allows reusing the same backward path across ensemble members rather than resampling for each model. In our experiments, we also mitigated cost by sampling only once per terminal state, but scaling to very large ensembles may still require additional engineering or approximations.
@@
Broader Impact. BGFNs provide a general mechanism for continued training and integration of multiple GFlowNets, opening the door to ensembles that combine different architectures, training regimes, or even loss functions. Although we did not explore these extensions in this work, the framework naturally supports them. This flexibility suggests that BGFNs could serve as a unifying tool for generative modeling in diverse domains, ranging from scientific discovery tasks to applications in reinforcement learning and probabilistic inference, provided appropriate care is taken in designing the ensemble and its reward structure.
@@
Supplementary Material
Appendix S1 Notation and objectives
@@
Let $R_{\theta}(x;\tau)\equiv\widehat{R}_{\theta}(x;\tau)$ be the current estimator. For a frozen ensemble $\mathcal{F}$, we define the marginal reference reward
$R_{\text{old}}(x)\;:=\;\sum_{k\in\mathcal{F}}\;\mathbb{E}_{\tau\sim P_{B}^{(k)}(\cdot\mid x)}\!\Big[\widehat{R}_{k}(x;\tau)\Big],$
i.e., a sum of member-wise marginals. Given $\alpha\in(0,1]$, we use
$\mathcal{L}_{\mathrm{boost}}(\tau)\;=\;\left(\log\frac{R_{\theta}(x;\tau)\;+\;\alpha\,R_{\phi}(x)}{\,R(x)\;-\;(1-\alpha)\,R_{\text{old}}(x)\,}\right)^{2}.$ (19)
Positivity of the denominator. To ensure $R(x)-(1-\alpha)R_{\text{old}}(x)>0$ uniformly, we replace $\alpha$ by a per-terminal state clamped value
$\alpha_{t}(x)\;=\;\mathrm{clip}\!\Big(\alpha,\ \alpha_{\min}(x),\ 1\Big),\qquad\alpha_{\min}(x)\;=\;\begin{cases}0,&R_{\text{old}}(x)=0,\\[4.0pt] 1-\dfrac{R(x)-\delta}{R_{\text{old}}(x)},&R_{\text{old}}(x)>0,\end{cases}$
@@
$R(x)\;-\;(1-\alpha_{t}(x))\,R_{\text{old}}(x)\ \geq\ \delta\ >0$
Special cases. If $R_{\text{old}}\equiv 0$, then $\mathcal{L}_{\mathrm{boost}}(\tau)=(\log R_{\theta}-\log R)^{2}$, i.e., it reduces to TB. If $R_{\text{old}}(x)=R(x)$, then $\mathcal{L}_{\mathrm{boost}}(\tau)=\big(\log(R_{\theta}/(\alpha R)+1)\big)^{2}$. (See Section S4 for the accompanying statements and proofs.)
@@
Estimating $\widehat{R}_{\theta}$ via importance sampling
@@
$\widehat{R}_{\theta}(x)\;\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\;Z_{\theta}\sum_{\tau:\,\tau\to x}P_{F}(\tau)=Z_{\theta}\,P_{F}(x).$ (20)
@@
$\widehat{R}_{\theta}(x)=\sum_{\tau:\,\tau\to x}Z_{\theta}P_{F}(\tau)\frac{P_{B}(\tau\mid x)}{P_{B}(\tau\mid x)}=\mathbb{E}_{\tau\sim P_{B}(\cdot\mid x)}\!\Big[\tfrac{Z_{\theta}\,P_{F}(\tau)}{P_{B}(\tau\mid x)}\Big].$ (21)
A Monte Carlo estimator with $K$ reverse paths is
$\widehat{R}_{K}(x)=\frac{1}{K}\sum_{i=1}^{K}\frac{Z_{\theta}\,P_{F}(\tau_{i})}{P_{B}(\tau_{i}\mid x)},\qquad\tau_{i}\sim P_{B}(\cdot\mid x).$ (22)
For a frozen ensemble, estimate each member marginal $\widehat{R}_{k}(x)\approx\frac{1}{K}\sum_{i}\frac{Z_{k}P_{F}^{(k)}(\tau_{i})}{P_{B}^{(k)}(\tau_{i}\mid x)}$ with $\tau_{i}\sim P_{B}^{(k)}(\cdot\mid x)$, and form $R_{\text{old}}(x)=\sum_{k}\widehat{R}_{k}(x)$. All computations are done in log-space (log-sum-exp) for numerical stability.
Appendix S2 Grid Environment:
@@ A Monte Carlo estimator with $K$ reverse paths is
Monte-Carlo marginalization over backward paths.
For terminal $x$, we approximate each member marginal with only one sample:
$\widehat{R}_{k}(x)\;\approx\;\frac{Z_{k}\,P_{F}^{(k)}(\tau)}{P_{B}^{(k)}(\tau\mid x)},\qquad\tau\sim P_{B}^{(k)}(\cdot\mid x),$
and then sum $\widehat{R}_{\text{old}}=\sum_{k}\widehat{R}_{k}(x)$.
@@
Baseline (K=1K{=}1): single forward/backward pair $(\pi_{F},\pi_{B})$ and a scalar $\log Z$. Boosting: freeze previously trained members and add a new forward/backward pair trained with the $\alpha$-Boosted objective (Sec. S1).
@@
$\widehat{R}_{k}(x)\;:=\;\frac{1}{B}\sum_{b=1}^{B}\frac{Z_{k}\,P_{F}^{(k)}(\tau^{(b)})}{P_{B}^{(k)}(\tau^{(b)}\mid x)}$
then aggregate $\widehat{R}(x)=\sum_{k}\widehat{R}_{k}(x)$. We report
@@
Appendix S3 Peptide Generation
@@
$\widehat{R}_{\text{old}}(y)\;=\;\sum_{k\in\mathcal{F}}Z_{k}\,P_{F}^{(k)}(y)\;=\;\sum_{k\in\mathcal{F}}\widehat{R}_{k}(y).$
@@
We train with AdamW on batches of on-policy rollouts. The objective is the TB regression in Eq. (18) (or the $\alpha$-Boosted variant in Eq. (19)). Baseline (K=1K{=}1) uses a single forward policy $\pi_{\theta}$ and a scalar $\log Z$; the backward policy is deterministic (pop-last-token) and not learned. Boosting: we freeze previously trained members and add a new forward policy and scalar trained with the $\alpha$-Boosted objective; the reference $\widehat{R}_{\text{old}}$ is the sum of terminal flows of the frozen members (Sec. S1).
@@
Appendix S4 Theorems and Proofs
@@
and, for a given $\alpha\in(0,1]$ and a marginal reference $\widehat{R}_{\text{old}}(x)\geq 0$, consider the per-trajectory boosted loss
$L_{\text{boost}}(\theta;\tau,x):=\Big(\log\big[\widehat{R}_{\theta}(x;\tau)+\alpha\,\widehat{R}_{\text{old}}(x)\big]-\log\big[R(x)-(1-\alpha)\,\widehat{R}_{\text{old}}(x)\big]\Big)^{2}.$ (24)
@@
    (no degradation) If $\widehat{R}_{\text{old}}(x)=R(x)$, then every stationary point of $L_{\text{boost}}$ (with respect to $\theta$) satisfies $Z_{\theta}=0$.
@@
    (residual focus) If $\widehat{R}_{\text{old}}(x)=0$, then $L_{\text{boost}}(\theta;\tau,x)$ coincides with the Trajectory Balance (TB) per-trajectory loss at $x$, and thus the stationary points of $L_{\text{boost}}$ are exactly the stationary points of the TB loss.
@@
(ii) Residual focus. If $\widehat{R}_{\text{old}}(x)=0$, then from (24)
@@
(i) No degradation. If $\widehat{R}_{\text{old}}(x)=R(x)$, then
$L_{\text{boost}}(\theta;\tau,x)=\Big(\log\!\big[1+\widehat{R}_{\theta}(x;\tau)/(\alpha R(x))\big]\Big)^{2}.$
@@
Given $N$ stages of GFlowNets $\{(Z_{i},P_{i,F},P_{i,B})\}_{i=1}^{N}$, define the estimated probability mass:
@@
where $\Gamma_{x}$ is the set of paths that terminate at $x$. For a target reward $R(x)$, if $\mathbb{E}_{\tau\sim Q}[L_{\text{boost}}]=0$, where $Q$ has full support over $\Gamma$, then $\hat{p}(x)\propto R(x)$.
Proof:.
For each stage $i$, set
$P_{i}(x):=\sum_{\tau\in\Gamma_{x}}P_{i,F}(\tau),\qquad S_{i}:=\{x:\ P_{i}(x)>0\}.$
The zero-boosting-loss condition with full-support $Q$ implies TB path-balance,
$Z_{N}\frac{P_{N,F}(\tau)}{P_{N,B}(\tau\mid x)}+R(x)\mathbf{1}_{\{x\notin S_{N}\}}=R(x)$ (26)
$\Rightarrow$ $Z_{N}\frac{P_{N,F}(\tau)}{P_{N,B}(\tau\mid x)}=R(x)\mathbf{1}_{\{x\in S_{N}\}}$
$\Rightarrow$ $\sum_{\tau\in\Gamma_{x}}Z_{N}P_{N,F}(\tau)=\sum_{\tau\in\Gamma_{x}}P_{N,B}(\tau\mid x)R(x)\mathbf{1}_{\{x\in S_{N}\}}$
$\Rightarrow$ $Z_{N}P_{N}(x)=R(x)\mathbf{1}_{\{x\in S_{N}\}}$
On the N-k iteration of $L_{\text{boost}}$ we had the TB condition valid for all trajectories expect on $\cup_{t=N-k+1}^{N}S_{t}$:
$Z_{N-k}\frac{P_{N-k,F}(\tau)}{P_{N-k,B}(\tau\mid x)}+R(x)\mathbf{1}_{\{x\notin S_{N-k}\}}=R(x);\quad\forall\tau\notin\cup_{t=N-k+1}^{N}S_{t}$ (27)
$\Rightarrow$ $Z_{N-k}P_{N-k}(x)=R(x)\mathbf{1}_{\{x\in S_{N-k}\}};\quad\forall\tau\notin\cup_{t=N-k+1}^{N}S_{t}$
Summing over all iteration of boosters we have:
$\sum_{k=1}^{N}Z_{k}P_{k}(x)=\ \sum_{k=1}^{N}R(x)\mathbf{1}_{\{x\in S_{k}\}}$ (28)
$\Rightarrow$ $\frac{\sum_{k=1}^{N}Z_{k}P_{k}(x)}{\sum_{x}R(x)}=\ \frac{R(x)}{\sum_{x}R(x)}$
Hence:
$\sum_{i=1}^{N}\frac{Z_{i}}{\sum_{j=1}^{N}Z_{j}}\sum_{\tau\in\Gamma_{x}}P_{i,F}(\tau)=\frac{R(x)}{Z}$

### v2

@@ 5.1 Synthetic Multimodal Environments
Training protocol.
We organized the experiments into three settings: (i) a single GFlowNet trained for 10,000 epochs (SINGLE), (ii) a baseline trained for 3,000 epochs followed by a first booster trained for an additional 7,000 epochs (BGFN-2), and (iii) the same baseline and first booster up to 6,000 epochs, after which a second booster is introduced and trained for 4,000 epochs (BGFN-3). To control for stochasticity, all models were trained with identical random seeds across conditions.
5.1.1 Results: BGFNs Overcome Exploration Bottlenecks and are Robust
Our results on the synthetic environments show a clear and consistent pattern, visualized in Figure 1 Figure 2 and Table 1:
@@ 5.2 Antimicrobial Peptide (AMP) Discovery
We evaluate BGFNs on a real combinatorial design task: the generation of antimicrobial peptides (AMPs) Trabucco et al. (2022); Pirtskhalava et al. (2021). Unlike prior work that fixes sequence length, here peptides have variable length from 1 to 10 amino acids. We train a Random Forest proxy to predict the probability of antimicrobial activity and define the GFlowNet reward as the logit of this probability. To avoid overweighting a small set of very-high-scoring sequences, we apply a 94% cutoff and clip the log-reward at zero for sequences above this threshold (i.e., all peptides with predicted activity $\geq 0.94$ share the same terminal reward). This choice preserves ranking signal below the threshold while mitigating oversampling to a few modes above it.
Training protocol (AMPs).
All models are trained for 3,000 epochs. We activate Booster 1 at epoch 1,200 and Booster 2 at epoch 2,400, keeping earlier stages frozen thereafter. We evaluate both on-policy and off-policy training under estimator noise levels $\varepsilon\in\{0.1,0.2,0.3\}$. In addition, we compare two loss instantiations for incorporating past rewards: the flow-additive and the target-residual variants. As in the synthetic setup, we reuse the same random seeds across conditions to ensure that observed differences arise from the training regime rather than luck.
5.2.1 Results: BGFNs Discover Substantially More Unique High-Reward Peptides
@@ Conclusion.
We introduced Boosted GFlowNets (BGFNs), a framework for sequentially training ensembles of GFlowNets. Our results demonstrate that BGFNs are safe to use: new boosters can be added to improve exploration without degrading the performance of previously trained components. This property provides a principled way of enhancing exploration in challenging environments, where the topology of the state space creates strong biases towards easy-to-reach modes, even under off-policy training. Across synthetic multimodal benchmarks and antimicrobial peptide discovery, BGFNs consistently improved mode coverage and robustness, validating the framework as a reliable extension of standard GFlowNet training.
Limitations.
@@
The main limitation of BGFNs is computational: each additional booster requires sampling from the backwards policy of all previously trained models in the ensemble. This increases training time on the order of $\mathcal{O}(KN)$, where $K$ denotes the number of trajectories sampled for a terminal node and $N$ the number of GFlowNets in the ensemble. An important exception is sequence environments with a unique backward path, in which the forward pass specifies the single valid backward trajectory; in practice, this allows reusing the same backward path across ensemble members rather than resampling for each model. In our experiments, we also mitigated cost by sampling only once per terminal state, but scaling to very large ensembles may still require additional engineering or approximations.
Broader Impact.
@@
BGFNs provide a general mechanism for continued training and integration of multiple GFlowNets, opening the door to ensembles that combine different architectures, training regimes, or even loss functions. Although we did not explore these extensions in this work, the framework naturally supports them. This flexibility suggests that BGFNs could serve as a unifying tool for generative modeling in diverse domains, ranging from scientific discovery tasks to applications in reinforcement learning and probabilistic inference, provided appropriate care is taken in designing the ensemble and its reward structure.
Acknowledgements
We acknowledge the support by the Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) (SEI-260003/020348/2025, SEI-260003/020694/2025) and the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) (404336/2023-0, 305692/2025-9).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] – Section 3 defines the formal objectives, balance conditions, and algorithms.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes] – The Limitations section discusses computational cost $\mathcal{O}(KN)$ of the ensemble.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] – Code and instructions will be provided in the supplementary material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes] – Theorems explicitly state assumptions on support and continuity.
    2. (b)
      Complete proofs of all theoretical results. [Yes] – Proofs are included in Appendix B.
    3. (c)
      Clear explanations of any assumptions. [Yes] – Assumptions are explained in theorems and in the background section.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes] – Code, data, and instructions will be released in the supplementary material.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes] – Training protocols, epochs, seeds, and loss variants are described in Section 5 and Appendix.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes] – Results are reported as mean ± standard deviation over 5 seeds.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes] – All experiments were run on CPUs only.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes] – The AMP dataset and proxy model are cited.
    2. (b)
      The license information of the assets, if applicable. [Yes] – The AMP dataset is publicly available; license details will be clarified in the supplementary material.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Acknowledgements
Supplementary Materials
Appendix A Notation and objectives
@@
Let ${R}_{\theta}(x;\tau)$ be the current trainable estimator. For a frozen ensemble $\mathcal{F}$, we define the induced reward terminal estimate as
$\widehat{R}(x)\;:=\;\sum_{k\in\mathcal{F}}\;\mathbb{E}_{\tau\sim P_{B}^{(k)}(\cdot\mid x)}\!\Big[Z_{k}\frac{P_{F}^{(k)}(\tau)}{P_{B}^{(k)}(\tau\mid x)}\Big],$
Given $\alpha\in[0,1]$, we define the boosted loss
$\mathcal{L}_{\mathrm{boost}}^{(k)}(\tau)\;=\;\left(\log\frac{R_{\theta}(x;\tau)\;+\;\alpha\,\widehat{R}_{k}(x)}{\,R(x)\;-\;(1-\alpha)\,\widehat{R}_{k}(x)\,}\right)^{2}.$ (19)
Here, $\widehat{R}_{k}(x)$ is a $k$-sample Monte Carlo estimate of the frozen model’s induced terminal estimate $\widehat{R}(x)$. To simplify notation we omit k and assume only one Monte Carlo sample from each model in the ensemble.
Positivity of the denominator. To ensure $R(x)-(1-\alpha)\widehat{R}(x)>0$ uniformly, we replace $\alpha$ by a per-terminal state clamped value
$\alpha_{t}(x)\;=\;\mathrm{clip}\!\Big(\alpha,\ \alpha_{\min}(x),\ 1\Big),\qquad\alpha_{\min}(x)\;=\;\begin{cases}0,&\widehat{R}(x)=0,\\[4.0pt] 1-\dfrac{R(x)-\delta}{\widehat{R}(x)},&\widehat{R}(x)>0,\end{cases}$
@@
$R(x)\;-\;(1-\alpha_{t}(x))\,\widehat{R}(x)\ \geq\ \delta\ >0$
Special cases. If $\widehat{R}\equiv 0$, then $\mathcal{L}_{\mathrm{boost}}(\tau)=(\log R_{\theta}-\log R)^{2}$, i.e., it reduces to TB. If $\widehat{R}(x)=R(x)$, then $\mathcal{L}_{\mathrm{boost}}(\tau)=\big(\log(R_{\theta}/(\alpha R)+1)\big)^{2}$. (See Section D for the accompanying statements and proofs.)
@@
Estimating $\widehat{R}$ via importance sampling
@@
$\widehat{R}(x)\;\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\;Z\sum_{\tau:\,\tau\to x}P_{F}(\tau)=Z\,P_{F}(x).$ (20)
@@
$\widehat{R}(x)=\sum_{\tau:\,\tau\to x}ZP_{F}(\tau)\frac{P_{B}(\tau\mid x)}{P_{B}(\tau\mid x)}=\mathbb{E}_{\tau\sim P_{B}(\cdot\mid x)}\!\Big[\tfrac{Z\,P_{F}(\tau)}{P_{B}(\tau\mid x)}\Big].$ (21)
Appendix B Grid Environment:
@@
Baseline (K=1K{=}1): single forward/backward pair $(\pi_{F},\pi_{B})$ and a scalar $\log Z$. Boosting: freeze previously trained members and add a new forward/backward pair trained with the $\alpha$-Boosted objective (Sec. A).
@@
$\widehat{R}^{k}(x)\;:=\;\frac{1}{B}\sum_{b=1}^{B}\frac{Z_{k}\,P_{F}^{(k)}(\tau^{(b)})}{P_{B}^{(k)}(\tau^{(b)}\mid x)}$
then aggregate $\widehat{R}(x)=\sum_{k}\widehat{R}^{k}(x)$. We report
@@
Appendix C Peptide Generation
@@
$\widehat{R}(y)\;=\;\sum_{k\in\mathcal{F}}Z_{k}\,P_{F}^{(k)}(y)\;=\;\sum_{k\in\mathcal{F}}\widehat{R}^{k}(y).$
@@
We train with AdamW on batches of on-policy rollouts. The objective is the TB regression in Eq. (18) (or the $\alpha$-Boosted variant in Eq. (19)). Baseline (K=1K{=}1) uses a single forward policy $\pi_{\theta}$ and a scalar $\log Z$; the backward policy is deterministic (pop-last-token) and not learned. Boosting: we freeze previously trained members and add a new forward policy and scalar trained with the $\alpha$-Boosted objective; the reference $\widehat{R}_{\text{old}}$ is the sum of terminal flows of the frozen members (Sec. A).
@@
Appendix D Theorems and Proofs
@@
and, for a given $\alpha\in(0,1]$ and a learned reward $\widehat{R}(x)\geq 0$, consider the per-trajectory boosted loss
$L_{\text{boost}}(\theta;\tau,x):=\Big(\log\big[\widehat{R}_{\theta}(x;\tau)+\alpha\,\widehat{R}(x)\big]-\log\big[R(x)-(1-\alpha)\,\widehat{R}(x)\big]\Big)^{2}.$ (23)
@@
    (no degradation) If $\widehat{R}(x)=R(x)$, then every stationary point of $L_{\text{boost}}$ (with respect to $\theta$) satisfies $Z_{\theta}=0$.
@@
    (residual focus) If $\widehat{R}(x)=0$, then $L_{\text{boost}}(\theta;\tau,x)$ coincides with the Trajectory Balance (TB) per-trajectory loss at $x$, and thus the stationary points of $L_{\text{boost}}$ are exactly the stationary points of the TB loss.
@@
(ii) Residual focus. If $\widehat{R}(x)=0$, then from (23)
@@
(i) No degradation. If $\widehat{R}(x)=R(x)$, then
$L_{\text{boost}}(\theta;\tau,x)=\Big(\log\!\big[1+\widehat{R}_{\theta}(x;\tau)/\alpha R(x)\big]\Big)^{2}.$
@@
Given $N$ stages of GFlowNets $\{(Z_{i},P_{i,F},P_{i,B})\}_{i=1}^{N}$, define
@@
where $\Gamma_{x}$ denotes the set of trajectories that terminate at $x$. For a target reward $R(x)$, if $\mathbb{E}_{\tau\sim Q}[L_{\text{boost}}]=0$ for a distribution $Q$ having full support over $\Gamma$, then $\hat{p}(x)\propto R(x)$.
Proof.
Let $\mathcal{X}$ be the set of terminal states and $\Gamma$ the set of complete trajectories. For $x\in\mathcal{X}$, let $\Gamma_{x}\subseteq\Gamma$ be the set of trajectories terminating at $x$, and for $S\subseteq\mathcal{X}$ define $\Gamma_{S}:=\bigcup_{x\in S}\Gamma_{x}$. For stage $i$, define the terminal marginal
$P_{i}(x)\;:=\;\sum_{\tau\in\Gamma_{x}}P_{i,F}(\tau),\qquad x\in\mathcal{X}.$
Because $Q$ has full support over $\Gamma$, we therefore have $L_{\text{boost}}(\tau)=0$ for all $\tau\in\Gamma$, and hence the defining boosted equality holds on every trajectory. Let us also consider an ordering of subsets $S_{1},\dots,S_{N}\subseteq\mathcal{X}$ and write
$U_{0}:=\varnothing,\qquad U_{i}:=\bigcup_{j=1}^{i}S_{j}.$
Stage 1 (classical TB on $S_{1}$). When the TB loss is zero on all trajectories terminating in $S_{1}$, the trajectory-balance identity holds; that is, for every $x\in S_{1}$ and every $\tau\in\Gamma_{x}$,
$Z_{1}\,P_{1,F}(\tau)\;=\;R(x)\,P_{1,B}(\tau\mid x).$ (25)
Summing (25) over $\tau\in\Gamma_{x}$ and using $\sum_{\tau\in\Gamma_{x}}P_{1,B}(\tau\mid x)=1$ yields
$Z_{1}\,P_{1}(x)\;=\;R(x)\qquad\forall x\in S_{1}.$ (26)
Thus the unnormalized terminal flow contributed by stage $1$ equals $R(x)$ on $S_{1}$.
Inductive claim. Assume that after stages $1,\dots,i-1$ the frozen ensemble explains exactly the terminals in $U_{i-1}$, namely
$\widehat{R}_{i-1}(x)\;=\;R(x)\,\mathbf{1}_{U_{i-1}}(x).$ (27)
We show that stage $i$ contributes unnormalized terminal flow equal to $R(x)$ on the newly covered terminals $S_{i}\setminus U_{i-1}$ and 0 on $S_{i}\cap U_{i-1}$.
Let $\alpha\in[0,1]$ denote the mixing parameter of the boosted loss, and define $\widehat{R}_{i}(x;\tau):=Z_{i}\,P_{i,F}(\tau)/P_{i,B}(\tau\mid x)$. Zero boosted loss implies that for every $x\in S_{i}$ and every $\tau\in\Gamma_{x}$,
$\widehat{R}_{i}(x;\tau)+\alpha\,\widehat{R}_{i-1}(x)\;=\;R(x)-(1-\alpha)\,\widehat{R}_{i-1}(x).$ (28)
If $x\in S_{i}\setminus U_{i-1}$, then $\widehat{R}_{i-1}(x)=0$ by (27), and (28) reduces to $\widehat{R}_{i}(x;\tau)=R(x)$ for all $\tau\in\Gamma_{x}$. Summing over $\Gamma_{x}$ gives
$Z_{i}\,P_{i}(x)\;=\;R(x)\qquad\forall x\in S_{i}\setminus U_{i-1}.$ (29)
If instead $x\in S_{i}\cap U_{i-1}$, then $\widehat{R}_{i-1}(x)=R(x)$ and (28) gives $\widehat{R}_{i}(x;\tau)=0$ for all $\tau\in\Gamma_{x}$, hence
$Z_{i}\,P_{i}(x)\;=\;0\qquad\forall x\in S_{i}\cap U_{i-1}.$ (30)
Therefore, for all $x\in\mathcal{X}$,
$Z_{i}\,P_{i}(x)\;=\;R(x)\,\mathbf{1}_{S_{i}\setminus U_{i-1}}(x).$ (31)
In particular, adding stage $i$ extends the explained set from $U_{i-1}$ to $U_{i}$ and preserves the form (27) with $i$ in place of i1i-1, completing the inductive step.
Summing (31) over $i=1,\dots,N$ then gives, for every $x\in\mathcal{X}$,
$\sum_{i=1}^{N}Z_{i}\,P_{i}(x)\;=\;R(x)\sum_{i=1}^{N}\mathbf{1}_{S_{i}\setminus U_{i-1}}(x)\;=\;R(x)\,\mathbf{1}_{U_{N}}(x).$ (32)
In particular, if $U_{N}=\mathcal{X}$, we have $\sum_{i=1}^{N}Z_{i}\,P_{i}(x)=R(x)$ for all $x\in\mathcal{X}$.
Finally, consider the sampling procedure that first draws an index $I$ with $P(I=i)=Z_{i}/\sum_{j=1}^{N}Z_{j}$ and then samples $x\sim P_{i}(\cdot)$ by running the stage-$i$ forward policy. The induced terminal distribution is
$\hat{p}(x)=\sum_{i=1}^{N}\frac{Z_{i}}{\sum_{j=1}^{N}Z_{j}}P_{i}(x)=\frac{1}{\sum_{j=1}^{N}Z_{j}}\sum_{i=1}^{N}Z_{i}\,P_{i}(x).$
Applying (32) with $U_{N}=\mathcal{X}$ yields
$\hat{p}(x)=\frac{R(x)}{\sum_{y\in\mathcal{X}}R(y)},$
and therefore $\hat{p}(x)\propto R(x)$, as claimed. ∎